A student takes out two loans, totaling $4000. One loan is at 6% interest, the other at 5% interest. the interest for one year is $215. what is the amount of each loan?
Answer by mananth(16946)

Part I 6.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + y= 4000 -------------1
Interest on both investments
6.00% x + 5.00% y= 215
Multiply by 100
6 x + 5 y= 21500.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -24000.00
Add this to (2)
0 x -1 y= -2500
divide by -1
y = 2500
Part I 6.00% $ 1500
Part II 5.00% $ 2500

CHECK
1500 --------- 6.00% ------- 90.00
2500 ------------- 5.00% ------- 125.00
Total -------------------- 215.00